[Brahms-l] Shift report 20040227 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Fri Feb 27 2004 - 23:58:10 EST
------------------------------------------------------------
Date: Fri Feb 27 23:58:10 2004, Shift: 16:00-24:00
Supervisor: rw, with: 
Log book#15 pages 68 - 68

Shift Summary:
------------------------------------------------------------
1. Beam is on at 18:00 till 22:30. The second store is made at 23:30.

2.  run    trg2     trg3
   10639             133k   trig2 is lost
   10640     33k     105k   trig2 is back
   10641    special request run FS 1/4A @ 40deg. Trig 6(100Hz)+ Trg 4,5(10Hz)
            1hr run
   10642    special run 2. Same FS field. Trg 4+5 (100Hz) 
   10643    TPC pedestal run
   10644    Non-TPC pedestal run
   
   10645    running  FS 1/2A at 8deg. MRS B2255 at 40 deg. 
   
3. No detector problem.

4. In the beginning of the run (10639) trg2 was not fired. It seems a connector
contact problem. After checking the trigger signals, the trigger came back.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Feb 27 23:58:33 2004

This archive was generated by hypermail 2.1.8 : Fri Feb 27 2004 - 23:58:53 EST